The perimeter of a rectangle is 22cm. The length is 5cm more than the width. Write and solve a system of equations to find the dimensions of the rectangle.
Answer by mananth(16946)


The perimeter of a rectangle is 22cm.
2(l+w)=22--------------------(1)

The length is 5cm more than the width.
l=w+5
substitute l in (1)
2(w+5+w)=22
2(2w+5)=22
4w+10=22
4w=12
w=3
if w=3 then l =8